OpenStudy (anonymous):

No idea where to start with this A 10cm-radius, 25-loop circular wire coil with 40 ohms of resistance sits in a 50T magnetic field. Over the next eight seconds, the field strength is reduced by one-fifth. a) How much voltage is induced in the coil? b) How much current flows in the coil while the field is being reduced? c) If the field strength is reduced during a two-second time interval instead, how much current would flow in the coil?

OpenStudy (irishboy123):

Faraday's law for induction \[V = - N \frac{d (BA)}{dt}\] for (i) N = 25, calc A(rea) from radius, A is fixed, so d(BA)/dt = A dB/dt (and a 1/5 reduction over 8 seconds gives dB = -1/5 * 50T, dt = 8).

OpenStudy (anonymous):

Must we assume plane of coil is perpendicular to magnetic field vector?

OpenStudy (irishboy123):

yes, good point flux ø = BA if perp and the law really relates dø/dt to emf. ø = 0 if parallel.....